OpenStudy (anonymous):

How do I graph x^2 / 4 + (y-3)^2 /4 =1

OpenStudy (anonymous):

its a circle centred at 0,3

OpenStudy (anonymous):

this is a circle with radius 2 and center at (0,3)

OpenStudy (anonymous):

oh okay, so it's points would be (2,3), (0,1), (-2,3), (0,5)

OpenStudy (anonymous):

\[\frac{x ^{2}}{4}+\frac{(y-3) ^{2}}{4}=1\rightarrow x ^{2}+(y-3)^{2}=4 \rightarrow x ^{2}+(y-3)^{2}=2^{2}\]